# textwidth=0.8textheight=0.9

### v1

@@ 6 Deep Operator KANs
This section introduces Deep Operator Kolmogorov–Arnold Networks (DeepOKANs), an emerging class of KANs paradigm designed to model entire operators, rather than pointwise input-output relationships. Instead of processing a finite-dimensional input vector, these architectures map an input function (e.g., boundary or initial conditions) to an output function (e.g., a PDE solution), thereby addressing the repeated-simulation bottleneck in parametric PDE settings, by learning a single surrogate that applies to any admissible input condition within a prescribed family. The section is organized into three parts. The first part describes the architecture and training process of DeepOKAN, focusing on its branch–trunk structure and the use of KANs to represent operators. It also explains how operator-level decomposition enables integration of physics-based constraints across the solution domain. The second part presents applications in areas such as solid mechanics, heat transfer, and electromagnetics, highlighting the model’s ability to approximate solution fields under varying conditions. The third part provides a comparative analysis between DeepOKAN and other operator learning methods, focusing on predictive accuracy and convergence behavior.
@@ 6.1 Architecture and Training Strategies
Similar to the DeepONet framework of Lu et al. (2021) [59], the DeepOKAN architecture is composed of two neural subnetworks as shown in Fig. 13: (i) a branch network that encodes the input function, represented via its evaluations at a set of fixed sensor locations $\bm{s}$, into a low-dimensional latent feature vector, and (ii) a trunk network that maps arbitrary spatiotemporal query points $\bm{\xi}=(\bm{x},t)$ to a corresponding set of latent basis functions. Unlike DeepONet, which typically employs MLPs in both subnetworks, DeepOKAN replaces them with KANs. These perform learnable univariate transformations along each input dimension, offering greater expressivity and flexibility in capturing localized, non-smooth, or highly nonlinear behaviors. Such a structure helps alleviate spectral bias commonly observed in standard operator learning [56, 234]. Formally, given an input function $v\in\mathcal{V}$, its discretized evaluation at sensor points is passed through the branch net, producing a feature vector $\bm{\beta}(v)=[\beta_{1},\beta_{2},\dots,\beta_{q}]^{\top}\in\mathbb{R}^{q}$. Simultaneously, for a query point $\bm{\xi}\in\Omega\times[0,T]$, the trunk network yields $\bm{\tau}(\bm{\xi})=[\tau_{1},\tau_{2},\dots,\tau_{q}]^{\top}\in\mathbb{R}^{q}$. The model prediction is obtained through an inner product,
$\mathcal{G}_{\bm{\theta}}(v)(\bm{\xi})=\sum_{k=1}^{q}\beta_{k}(v)\cdot\tau_{k}(\bm{\xi}),$ (21)
where $\bm{\theta}$ denotes all trainable parameters across both subnetworks. Equation (21) defines a nonlinear operator $\mathcal{G}_{\bm{\theta}}\>:\>\mathcal{V}\to\mathcal{A}$ that maps input functions to output solution functions over the spatiotemporal domain. To guide the training of DeepOKAN, a composite loss function is used that combines both data-driven and physics-based terms. Let $\{v^{(j)}\}_{j=1}^{N}$ denote a set of training input functions sampled from the input space $\mathcal{V}$, each associated with an output governed by the underlying PDE.
@@
Once trained, DeepOKAN generalizes across a wide class of input functions and provides continuous predictions over the spatiotemporal domain without requiring re-solving the PDE for new inputs. Its use of dimension-wise KAN layers reduces the complexity of learning high-dimensional operators and improves training stability, particularly in challenging physical domains such as fluid dynamics, wave propagation, and solid mechanics [56]. Furthermore, normalization of inputs and dynamic loss balancing strategies can be employed to further enhance numerical stability and convergence during training.
To better address the challenges of operator learning, various architectural variants have been proposed by modifying the network components within the branch and trunk subnetworks. For example, Kiyani et al. [22] introduced a simplified architecture in which the trunk network is replaced by a KAN, trained without incorporating any physics-based loss. In contrast, Abueidda et al. [56] presented DeepOKAN, a neural operator that utilizes Gaussian RBF KANs in both subnetworks, demonstrating enhanced performance and interpretability. Their formulation serves as a performance benchmark and opens avenues for exploring other KAN variants, such as multi-quadratic RBFs and B-splines. Wang et al. [115] proposed a hybrid architecture combining a FNO in the branch network with a KAN-based trunk, leveraging the complementary strengths of spectral and localized representations. Additionally, Yu et al. [116] developed DeepOHeat-v1, which integrates adaptive KANs with learnable activation functions in the trunk to capture complex multi-scale thermal patterns. This model also incorporates a separable training strategy and GMRES-based (Generalized Minimal Residual Method) refinement to improve computational efficiency and reliability. Expanding on the DeepONet formulation, Toscano et al. [117] proposed DeepOKKAN, which incorporates KKANs characterized by MLP-based inner blocks and radial basis function outer blocks, offering a structured and interpretable architecture for operator learning. Collectively, these variants underscore the flexibility of KAN-based operator networks and their increasing significance in physics-informed machine learning frameworks.
@@ 6.2 Applications
DeepOKAN Introduced neural operators to model intricate relationships between input parameters and output fields; Used Gaussian RBFs for both branch and trunk networks; Applied to 1D sinusoidal waves, 2D orthotropic elasticity, and the transient Poisson’s problem. [56]
@@
Hybrid Operators Employed an MLP branch and Chebyshev KAN trunk; Trained concurrently; Applied to predicting displacement and phase fields from crack size, boundary displacements in 1D bars, and notched structures under tensile/shear loading. [22]
@@
Operator KANs Analyzed the performance of KANs and MLPs in both shallow and deep network models; Used spline-based KANs; Applied to Burgers’ equation, 1D and 2D Darcy Flow, and an Elastic Plate problem. [75]
@@
EFKAN Employed FNO as a branch and KAN (B-spline, degree 3, grid 5) as a trunk network; Mapped resistivity models to apparent resistivity and phase at target locations and frequencies; Applied to MT forward modeling in the frequency domain. [115]
@@
DeepOHeat-v1 Expanded DeepOHeat with Chebyshev KAN trunk; Used separable training and hybrid GMRES-based refinement for efficiency and reliability; Applied to 2D power maps, 3D volumetric power distribution, and thermal optimization of floorplans. [116]
@@
DeepOKKAN Employed cKAN and KKAN variants with RBF-based trunk blocks within the DeepOKKAN framework; Enhanced with QR-DeepONet reparameterization for improved stability and accuracy; Applied to solving the 1D Burgers’ equation. [117]
@@
In the world of solid mechanics, Kiyani et al. [22] used DeepOKAN to predict displacement and phase fields directly from initial crack geometry and boundary conditions, achieving high accuracy in 1D and 2D settings. The model bypasses traditional solver limitations by learning the operator mapping, resulting in reduced computational cost and data requirements. As shown in Fig. 14(a), DeepOKAN captures steep gradients in damage and displacement fields with close agreement to reference solutions. Its ability to represent nonlinear multiscale features makes it a promising tool for efficient and accurate fracture prediction.
@@
Figure 14: (a) Predicted crack branching and corresponding absolute error obtained using DeepOKAN. The model is trained using phase-field data from 45 randomly selected initial notches and tested on 5 unseen samples; the branch network processes the phase-field inputs, while the trunk network encodes spatial coordinates; each subnetwork contains 7 hidden layers with 100 nodes per layer, with 300 neurons in the final branch layer and 900 in the trunk output layer. This panel is adopted from [22]. (b) Predicted solution and absolute error for the 2D orthotropic elasticity problem using DeepOKAN. The model, with a medium architectural complexity with approximately 7,200 trainable parameters, is adopted from [56].(c) Predicted and absolute error of apparent resistivity and phase by EFKAN for a smooth resistivity model. The network is trained on data from 22 frequencies and 22 spatial locations, and evaluated on a denser grid of 64 frequencies and 64 coordinates. This panel is adopted from [115].(d) Evaluation of predictive accuracy and refinement capability in operator learning. DeepOHeat-v1 employs a separable architecture with a shared branch network and three independent trunk networks based on Chebyshev KANs, while the hybrid method refines the neural predictions by correcting residuals using GMRES. This panel is adopted from [116].
@@
In the context of fast surrogate modeling for complex mechanical systems, Abueidda et al. [56] proposed DeepOKAN, a surrogate model based on KANs with Gaussian RBFs, enabling efficient operator learning across diverse mechanical applications. The model has been applied in both data-driven and physics-informed settings, including 1D wave propagation, 2D linear elasticity, and time-dependent Poisson equations. These examples cover a range of static and dynamic problems. As shown in Fig. 14(b), the predicted fields closely match reference solutions, indicating accurate performance across varying temporal and spatial domains. For geophysical electromagnetics, Wang et al. [115] introduced the EFKAN framework, a neural operator framework that incorporates KANs into the projection layers. This design replaces standard MLP components to improve flexibility and representation in complex geophysical settings. The model has been tested on smooth and heterogeneous resistivity distributions generated via spectral methods, showing accurate predictions of apparent resistivity and phase responses (see Fig. 14(c)). EFKAN offers an efficient surrogate for MT simulations governed by Maxwell’s equations, suitable for integration into data-driven electromagnetic inversion workflows. In the area of electronics and heat management, Yu et al. [116] adaptively capture multiscale thermal patterns using the DeepOHeat-v1 framework. DeepOHeat-v1 is highly applicable to thermally aware floor planning and optimization in solid-state electronic systems. Fig. 14(d) illustrates the refinement process for thermal field predictions, where initial operator network outputs exhibit notable discrepancies. The refinement corrects these inaccuracies by leveraging the error feedback, producing thermally consistent solutions aligned with the ground truth.
@@ 6.3 Comparison with Other Operators
1 D DeepONet SiLU Shallow 2.89 % 5.58 % 9.42 % [75]
@@
2 D DeepONet tanh 147000 1.62 % 2.25 % 3.47 % [66]
@@
Figure 15: Comparison of prediction accuracy between different operator learning frameworks, focusing on error behavior under varying architectural depths, complexity levels, and physical settings.(a) Absolute error maps highlighting the different network architectures on prediction accuracy in heterogeneous flow fields. This panel is adopted from [75].(b) Absolute error compared to the FEA ground-truth in a high-complexity case with 50600 learnable parameters for DeepOKAN and 51204 learnable parameters for DeepONet. This panel is adopted from [56].(c) Error using TeData-III containing 100 resistivity profiles with rectangular anomalies, simulated at 64 frequencies (0.005–12.589 Hz). This panel is adopted from [115].
@@ 6.3.1 Accuracy Comparison
Understanding how operator learning models behave across varying network depths is crucial for evaluating their expressive power and computational efficiency. Pant et al. [75] systematically studied the behavior of shallow and deep architectures using KAN and MLP-based models within operator learning frameworks. Table 7 presents a summary of their results on the 1D Darcy flow problem. In shallow configurations, the KAN-based DeepOKAN shows a significant reduction in error compared to the MLP-based DeepONet, improving accuracy by over 85%. In deeper settings, DeepOKAN also performs better, with an error reduction of approximately 65%, although the performance gap becomes narrower as both models benefit from increased depth. Furthermore, Fig. 15(a) illustrates the model outputs for the more complex 2D Darcy flow scenario. The visualizations reveal that DeepOKAN retains strong performance in shallow architectures, capturing key spatial patterns with fewer parameters. This supports the conclusion that shallow KAN models offer high expressiveness and efficiency, making them well-suited for physics-based learning tasks that demand accuracy without high computational overhead.
@@
One important factor influencing the accuracy of operator learning methods is the number of available data samples. In many practical scenarios, obtaining large amounts of data may not be feasible due to cost, time, or physical limitations. Therefore, a desirable model should maintain high predictive accuracy even when the number of samples is limited. In particular, some works such as [115] and [56] have reported the effect of sample size on model performance in specific examples. Abueidda et al. [56] reported the behavior of DeepOKAN and DeepONet under different numbers of data samples, providing useful insights into how sample count affects model performance. Fig. 15(b) presents the prediction errors of DeepONet and DeepOKAN for the Transient Poisson’s problem under different sample settings. The results show that while the performance of DeepONet degrades significantly with fewer samples, DeepOKAN maintains relatively stable accuracy.
@@
In real-world physics-based learning tasks, input data is often contaminated with noise, making it essential to assess the robustness of operator learning models under such conditions. Understanding how models respond to increasing noise levels provides insights into their generalization ability and reliability in practical scenarios. This aspect has been explored in studies [75, 66] through the solution of the Darcy flow problem under varying levels of input noise. Table 7 summarizes their results across different dimensions and model configurations. In the 1D case, as noise levels increase from 0% to 10%, DeepOKAN consistently maintains lower error compared to DeepONet. Specifically, in shallow architectures, DeepOKAN shows improvements of approximately 22% at 5% noise and around 9% at 10% noise. For deep models, the trend persists with DeepOKAN maintaining a smaller error, showing a roughly 7% advantage at 5% noise and 2% at 10% noise. These findings suggest that while both models experience degraded performance as noise increases, DeepOKAN demonstrates more robust behavior, especially in shallower settings. In the 2D Darcy flow problem, while DeepONet initially shows lower error under noise-free conditions, its performance degrades more sharply with added noise. In contrast, DeepOKAN exhibits stable accuracy across all tested noise levels, with only a slight increase in error, indicating better resilience to noisy inputs in this case.
@@
Exploring different architectural designs is crucial for enhancing the flexibility and expressiveness of operator learning models. One promising direction involves integrating KANs with advanced structures such as Fourier layers. For example, Wang et al. [115] proposed EFKAN, which combines KAN with FNO. Their results show that EFKAN achieves higher accuracy in estimating both apparent resistivity and phase. Fig. 15(c) illustrates that both EFNO and EFKAN can produce reasonable results for apparent resistivity and phase. However, on average, EFKAN reduces the relative $\ell_{1}$-norm error by around 20–25% across the different components, particularly improving phase accuracy without sacrificing performance on resistivity. Therefore, when integrated into more advanced operator frameworks like EFKAN, KANs are a compelling choice for tasks in computational physics.
@@
Figure 16: Convergence behavior of DeepOKAN and other operator networks across various problems. (a) Comparison of training loss for Burgers’ equation with fixed viscosity (0.01) and random initial conditions sampled from a Gaussian random field. The shallow network uses one hidden layer for both the branch and trunk, while the deep version uses 7 MLP layers and 4 KAN layers in each. This panel is reproduced from [75]. (b) Training loss curves for the transient Poisson equation with zero source and initial conditions. High-complexity DeepOKAN and DeepONet models have 50600 and 51204 parameters, respectively; low-complexity versions have 12650 and 13104 parameters. This panel is reproduced from [56]. (c) Relative $\mathcal{L}^{2}$ error convergence on the test set for Burgers’ equation with lower viscosity ($1/(100\pi)$). DeepOKKAN, DeepOcKAN, and DeepONet models have 155920, 107328, and 131700 parameters, respectively; QR-based models use over 46000 parameters. This panel is reproduced from [117]. (d) Training loss comparison of EFNO and EFKAN on the small-scale TrData dataset. This panel is reproduced from [115].

### v2

@@ 6 Deep-operator KANs
This section introduces Deep-operator Kolmogorov–Arnold Networks (DeepOKANs), an emerging class of KANs paradigm designed to model entire operators, rather than pointwise input-output relationships. Instead of processing a finite-dimensional input vector, these architectures map an input function (e.g., boundary or initial conditions) to an output function (e.g., a PDE solution), thereby addressing the repeated-simulation bottleneck in parametric PDE settings, by learning a single surrogate that applies to any admissible input condition within a prescribed family. The section is organized into three parts. The first part describes the architecture and training process of DeepOKAN, focusing on its branch–trunk structure and the use of KANs to represent operators. It also explains how operator-level decomposition enables integration of physics-based constraints across the solution domain. The second part presents applications in areas such as solid mechanics, heat transfer, and electromagnetics, highlighting the model’s ability to approximate solution fields under varying conditions. The third part provides a comparative analysis between DeepOKAN and other operator learning methods, focusing on predictive accuracy and convergence behavior.
@@ 6.1 Architecture and Training Strategies
Similar to the DeepONet framework of Lu et al. (2021) lu2021learning , the DeepOKAN architecture is composed of two neural subnetworks as shown in Fig. 14: (i) a branch network that encodes the input function, represented via its evaluations at a set of fixed sensor locations $\bm{s}$, into a low-dimensional latent feature vector, and (ii) a trunk network that maps arbitrary spatiotemporal query points $\bm{\xi}=(\bm{x},t)$ to a corresponding set of latent basis functions. Unlike DeepONet, which typically employs MLPs in both subnetworks, DeepOKAN replaces them with KANs. These perform learnable univariate transformations along each input dimension, offering greater expressivity and flexibility in capturing localized, non-smooth, or highly nonlinear behaviors. Such a structure helps alleviate spectral bias commonly observed in standard operator learning abueidda2025deepokan , Wang2020NTKPinns . Formally, given an input function $v\in\mathcal{V}$, its discretized evaluation at sensor points is passed through the branch network with parameters $\bm{\theta}_{b}$, producing a feature vector $\bm{\beta}(v;\bm{\theta}_{b})=[\beta_{1},\beta_{2},\dots,\beta_{q}]^{\top}\in\mathbb{R}^{q}$. Simultaneously, for a query point $\bm{\xi}\in\Omega\times[0,T]$, the trunk network with parameters $\bm{\theta}_{t}$ yields $\bm{\tau}(\bm{\xi};\bm{\theta}_{t})=[\tau_{1},\tau_{2},\dots,\tau_{q}]^{\top}\in\mathbb{R}^{q}$. The model prediction is obtained through an inner product,
$\mathcal{G}_{\bm{\theta}}(v)(\bm{\xi})=\sum_{k=1}^{q}\beta_{k}(v;\bm{\theta}_{b})\cdot\tau_{k}(\bm{\xi};\bm{\theta}_{t}),$ (21)
where $\bm{\theta}_{b}$ and $\bm{\theta}_{t}$ denote the parameter vectors of the branch and trunk networks, respectively, and the full parameter set is $\bm{\theta}=\{\bm{\theta}_{b},\bm{\theta}_{t}\}$. Equation (21) defines a nonlinear operator $\mathcal{G}_{\bm{\theta}}\>:\>\mathcal{V}\to\mathcal{A}$ that maps input functions to output solution functions over the spatiotemporal domain. To guide the training of DeepOKAN, a composite loss function is used that combines both data-driven and physics-based terms. Let $\{v^{(j)}\}_{j=1}^{N}$ denote a set of training input functions sampled from the input space $\mathcal{V}$, each associated with an output governed by the underlying PDE.
@@
Once trained, DeepOKAN generalizes across a wide class of input functions and provides continuous predictions over the spatiotemporal domain without requiring re-solving the PDE for new inputs. Its use of dimension-wise KAN layers reduces the complexity of learning high-dimensional operators and improves training stability, particularly in challenging physical domains such as fluid dynamics, wave propagation, and solid mechanics abueidda2025deepokan . Furthermore, normalization of inputs and dynamic loss balancing strategies can be employed to further enhance numerical stability and convergence during training.
To better address the challenges of operator learning, various architectural variants have been proposed by modifying the network components within the branch and trunk subnetworks. For example, Kiyani et al. kiyani2025predicting introduced a simplified architecture in which the trunk network is replaced by a KAN, trained without incorporating any physics-based loss. In contrast, Abueidda et al. abueidda2025deepokan presented DeepOKAN, a neural operator that utilizes Gaussian RBF KANs in both subnetworks, demonstrating enhanced performance and interpretability. Their formulation serves as a performance benchmark and opens avenues for exploring other KAN variants, such as multi-quadratic RBFs and B-splines. Wang et al. wang2025efkan proposed a hybrid architecture combining a FNO in the branch network with a KAN-based trunk, leveraging the complementary strengths of spectral and localized representations. Additionally, Yu et al. yu2025deepoheat developed DeepOHeat-v1, which integrates adaptive KANs with learnable activation functions in the trunk to capture complex multi-scale thermal patterns. This model also incorporates a separable training strategy and GMRES-based (Generalized Minimal Residual Method) refinement to improve computational efficiency and reliability. Expanding on the DeepONet formulation, Toscano et al. toscano2025kkans proposed DeepOKKAN, which incorporates KKANs characterized by MLP-based inner blocks and radial basis function outer blocks, offering a structured and interpretable architecture for operator learning. Collectively, these variants underscore the flexibility of KAN-based operator networks and their increasing significance in physics-informed machine learning frameworks.
@@ 6.2 Applications
DeepOKAN Introduced neural operators to model intricate relationships between input parameters and output fields; Used Gaussian RBFs for both branch and trunk networks; Applied to 1D sinusoidal waves, 2D orthotropic elasticity, and the transient Poisson’s problem. abueidda2025deepokan
@@
Hybrid Operators Employed an MLP branch and Chebyshev KAN trunk; Trained concurrently; Applied to predicting displacement and phase fields from crack size, boundary displacements in 1D bars, and notched structures under tensile/shear loading. kiyani2025predicting
@@
Operator KANs Analyzed the performance of KANs and MLPs in both shallow and deep network models; Used spline-based KANs; Applied to Burgers’ equation, 1D and 2D Darcy Flow, and an Elastic Plate problem. pant2025mlps
@@
EFKAN Employed FNO as a branch and KAN (B-spline, degree 3, grid 5) as a trunk network; Mapped resistivity models to apparent resistivity and phase at target locations and frequencies; Applied to MT forward modeling in the frequency domain. wang2025efkan
@@
DeepOHeat-v1 Expanded DeepOHeat with Chebyshev KAN trunk; Used separable training and hybrid GMRES-based refinement for efficiency and reliability; Applied to 2D power maps, 3D volumetric power distribution, and thermal optimization of floorplans. yu2025deepoheat
@@
DeepOKKAN Employed cKAN and KKAN variants with RBF-based trunk blocks within the DeepOKKAN framework; Enhanced with QR-DeepONet reparameterization for improved stability and accuracy; Applied to solving the 1D Burgers’ equation. toscano2025kkans
PO-CKAN Proposed a physics-informed deep-operator framework combining DeepONet with Chunk-rational KAN (CKAN) modules; Used CKAN for both branch and trunk nets; Applied to Burgers’, Eikonal, and diffusion-reaction equations. wu2025po
@@
In the world of solid mechanics, Kiyani et al. kiyani2025predicting used DeepOKAN to predict displacement and phase fields directly from initial crack geometry and boundary conditions, achieving high accuracy in 1D and 2D settings. The model bypasses traditional solver limitations by learning the operator mapping, resulting in reduced computational cost and data requirements. As shown in Fig. 15(a), DeepOKAN captures steep gradients in damage and displacement fields with close agreement to reference solutions. Its ability to represent nonlinear multiscale features makes it a promising tool for efficient and accurate fracture prediction.
@@
Figure 15: (a) Predicted crack branching and corresponding absolute error obtained using DeepOKAN. The model is trained using phase-field data from 45 randomly selected initial notches and tested on 5 unseen samples; the branch network processes the phase-field inputs, while the trunk network encodes spatial coordinates; each subnetwork contains 7 hidden layers with 100 nodes per layer, with 300 neurons in the final branch layer and 900 in the trunk output layer. This panel is adopted from kiyani2025predicting . (b) Predicted solution and absolute error for the 2D orthotropic elasticity problem using DeepOKAN. The model, with a medium architectural complexity with approximately 7,200 trainable parameters. This panel is adopted from abueidda2025deepokan . (c) Predicted and absolute error of apparent resistivity and phase by EFKAN for a smooth resistivity model. The network is trained on data from 22 frequencies and $2$ spatial locations, and evaluated on a denser grid of 64 frequencies and 64 coordinates. This panel is adopted from wang2025efkan .(d) Evaluation of predictive accuracy and refinement capability in operator learning. DeepOHeat-v1 employs a separable architecture with a shared branch network and three independent trunk networks based on Chebyshev KANs, while the hybrid method refines the neural predictions by correcting residuals using GMRES. This panel is adopted from yu2025deepoheat .
@@
In the context of fast surrogate modeling for complex mechanical systems, Abueidda et al. abueidda2025deepokan proposed DeepOKAN, a surrogate model based on KANs with Gaussian RBFs, enabling efficient operator learning across diverse mechanical applications. The model has been applied in both data-driven and physics-informed settings, including 1D wave propagation, 2D linear elasticity, and time-dependent Poisson equations. These examples cover a range of static and dynamic problems. As shown in Fig. 15(b), the predicted fields closely match reference solutions, indicating accurate performance across varying temporal and spatial domains. For geophysical electromagnetics, Wang et al. wang2025efkan introduced the EFKAN framework, a neural operator framework that incorporates KANs into the projection layers. This design replaces standard MLP components to improve flexibility and representation in complex geophysical settings. The model has been tested on smooth and heterogeneous resistivity distributions generated via spectral methods, showing accurate predictions of apparent resistivity and phase responses (see Fig. 15(c)). EFKAN offers an efficient surrogate for MT simulations governed by Maxwell’s equations, suitable for integration into data-driven electromagnetic inversion workflows. In the area of electronics and heat management, Yu et al. yu2025deepoheat adaptively capture multiscale thermal patterns using the DeepOHeat-v1 framework. DeepOHeat-v1 is highly applicable to thermally aware floor planning and optimization in solid-state electronic systems. Fig. 15(d) illustrates the refinement process for thermal field predictions, where initial operator network outputs exhibit notable discrepancies. The refinement corrects these inaccuracies by leveraging the error feedback, producing thermally consistent solutions aligned with the ground truth.
@@ 6.3 Comparison with Other Operators
1 D DeepONet SiLU Shallow 2.89 % 5.58 % 9.42 % pant2025mlps
@@
2 D DeepONet tanh 147000 1.62 % 2.25 % 3.47 % shukla2024comprehensive
@@
Figure 16: Comparison of prediction accuracy between different operator learning frameworks, focusing on error behavior under varying architectural depths, complexity levels, and physical settings.(a) Absolute error maps highlighting the different network architectures on prediction accuracy in heterogeneous flow fields. This panel is adapted from pant2025mlps . (b) Absolute error compared to the FEA ground-truth in a high-complexity case with 50600 learnable parameters for DeepOKAN and 51204 learnable parameters for DeepONet. This panel is adapted from abueidda2025deepokan . (c) Error using TeData-III containing 100 resistivity profiles with rectangular anomalies, simulated at 64 frequencies (0.005–12.589 Hz). This panel is adapted from wang2025efkan .
@@ 6.3.1 Accuracy Comparison
Understanding how operator learning models behave across varying network depths is crucial for evaluating their expressive power and computational efficiency. Pant et al. pant2025mlps systematically studied the behavior of shallow and deep architectures using KAN and MLP-based models within operator learning frameworks. Table 7 presents a summary of their results on the 1D Darcy flow problem. In shallow configurations, the KAN-based DeepOKAN shows a significant reduction in error compared to the MLP-based DeepONet, improving accuracy by over 85%. In deeper settings, DeepOKAN also performs better, with an error reduction of approximately 65%, although the performance gap becomes narrower as both models benefit from increased depth. Furthermore, Fig. 16(a) illustrates the model outputs for the more complex 2D Darcy flow scenario. The visualizations reveal that DeepOKAN retains strong performance in shallow architectures, capturing key spatial patterns with fewer parameters. This supports the conclusion that shallow KAN models offer high expressiveness and efficiency, making them well-suited for physics-based learning tasks that demand accuracy without high computational overhead.
@@
One important factor influencing the accuracy of operator learning methods is the number of available data samples. In many practical scenarios, obtaining large amounts of data may not be feasible due to cost, time, or physical limitations. Therefore, a desirable model should maintain high predictive accuracy even when the number of samples is limited. In particular, some works such as wang2025efkan and abueidda2025deepokan have reported the effect of sample size on model performance in specific examples. Abueidda et al. abueidda2025deepokan reported the behavior of DeepOKAN and DeepONet under different numbers of data samples, providing useful insights into how sample count affects model performance. Fig. 16(b) presents the prediction errors of DeepONet and DeepOKAN for the Transient Poisson’s problem under different sample settings. The results show that while the performance of DeepONet degrades significantly with fewer samples, DeepOKAN maintains relatively stable accuracy.
@@
In real-world physics-based learning tasks, input data is often contaminated with noise, making it essential to assess the robustness of operator learning models under such conditions. Understanding how models respond to increasing noise levels provides insights into their generalization ability and reliability in practical scenarios. This aspect has been explored in studies pant2025mlps , shukla2024comprehensive through the solution of the Darcy flow problem under varying levels of input noise. Table 7 summarizes their results across different dimensions and model configurations. In the 1D case, as noise levels increase from 0% to 10%, DeepOKAN consistently maintains lower error compared to DeepONet. Specifically, in shallow architectures, DeepOKAN shows improvements of approximately 22% at 5% noise and around 9% at 10% noise. For deep models, the trend persists with DeepOKAN maintaining a smaller error, showing a roughly 7% advantage at 5% noise and 2% at 10% noise. These findings suggest that while both models experience degraded performance as noise increases, DeepOKAN demonstrates more robust behavior, especially in shallower settings. In the 2D Darcy flow problem, while DeepONet initially shows lower error under noise-free conditions, its performance degrades more sharply with added noise. In contrast, DeepOKAN exhibits stable accuracy across all tested noise levels, with only a slight increase in error, indicating better resilience to noisy inputs in this case.
@@
Exploring different architectural designs is crucial for enhancing the flexibility and expressiveness of operator learning models. One promising direction involves integrating KANs with advanced structures such as Fourier layers. For example, Wang et al. wang2025efkan proposed EFKAN, which combines KAN with FNO. Their results show that EFKAN achieves higher accuracy in estimating both apparent resistivity and phase. Fig. 16(c) illustrates that both EFNO and EFKAN can produce reasonable results for apparent resistivity and phase. However, on average, EFKAN reduces the relative $\ell_{1}$-norm error by around 20–25% across the different components, particularly improving phase accuracy without sacrificing performance on resistivity. Therefore, when integrated into more advanced operator frameworks like EFKAN, KANs are a compelling choice for tasks in computational physics.
@@
Figure 17: Convergence behavior of DeepOKAN and other operator networks across various problems. (a) Comparison of training loss for Burgers’ equation with fixed viscosity (0.01) and random initial conditions sampled from a Gaussian random field. The shallow network uses one hidden layer for both the branch and trunk, while the deep version uses 7 MLP layers and 4 KAN layers in each. This panel is reproduced from pant2025mlps . (b) Training loss curves for the transient Poisson equation with zero source and initial conditions. High-complexity DeepOKAN and DeepONet models have 50600 and 51204 parameters, respectively; low-complexity versions have 12650 and 13104 parameters. This panel is reproduced from abueidda2025deepokan . (c) Relative $\mathcal{L}^{2}$ error convergence on the test set for Burgers’ equation with lower viscosity ($1/(100\pi)$). DeepOKKAN, DeepOcKAN, and DeepONet models have 155920, 107328, and 131700 parameters, respectively; QR-based models use over 46000 parameters. This panel is reproduced from toscano2025kkans . (d) Training loss comparison of EFNO and EFKAN on the small-scale TrData dataset. This panel is reproduced from wang2025efkan .